OpenStudy (anonymous):

a 25 foot wire is to be cut so that the longer piece is one foot longer than 5 times the shorter piece. Find the length of each piece

OpenStudy (hoblos):

x= length of longer piece y= length of shorter piece x+y=25 x=5y+1 replace this equation in the first eq 5y+1+y=25 y=4 x=21

OpenStudy (anonymous):

5x+1+x=25 where x is shorter lenght

OpenStudy (anonymous):

x=4 therefore the longer length is 5x+1 which is 4*4+1=21